OpenStudy (boxman61):

solve using one-to-one property ln(x+3)=ln(x+1)=ln8

OpenStudy (anonymous):

wrong!!!

OpenStudy (boxman61):

the given answer is 1 but i cannot work it out to equal 1...\[\ln (x+3)+\ln (+1)=\ln 8\] \[(x+3)+(x+1)=8\]\[2x+4=8\]\[2x=4\]\[x=2\]

OpenStudy (boxman61):

thats how i work it out... book says answer is 1

OpenStudy (anonymous):